Thirty ounces of pure silver are added to 50oz of a silver alloy that is 20% silver. What is the percent concentration of the resulting alloy?
Answer by RAY100(1637)

100% * 30 + 20% * 50 = (x%) * (30+50)
.
1*30 +.2*50 = x* 80
.
30+10= 80x
.
40=80x
.
.5 = x,,,,,,,or 50%
.
.
check,,,,30 + 10 = .5*80=40,,,,,ok
.